Three very large parallel non-conducting planes (i
Answer #1

A) E = sigma/(2*epsilon_0)

sigma is the surface charge density

B) at A

E = [sigma/(2*epsilon_o)] + [sigma/(2*epsilon_o)] =

E = [sigma/(epsilon_o)].........

at B

E = [sigma/(2*epsilon_o)] - [sigma/(2*epsilon_o)]   + [3*sigma/(2*epsilon_o)] = [3*sigma/(2*epsilon_o)]


at P


E = [sigma/(2*epsilon_o)] - [sigma/(2*epsilon_o)]   - [3*sigma/(2*epsilon_o)] = [-3*sigma/(2*epsilon_o)]

C) at B

Force on Q is F = Q*E = Q*[3*sigma/(2*epsilon_o)] = 3*Q*sigma/(2*epsilon_0)


at A

force on Q is F = Q*E = Q*sigma/(epsilon_o)